# K-theory and localizing invariants of large categories

### v2

@@ Appendix B Essential image of a homological epimorphism
Consider a homological epimorphism where and are small stable idempotent-complete categories. We know that if the category is compactly generated, then In particular, the essential image of is a stable subcategory of In this case by Thomason’s theorem every object of of the form can be lifted to
It turns out that if is not compactly generated then the image of might not be a stable subcategory of More precisely, the following result shows that in general for a dualizable category there may exist an object such that the object cannot be lifted to
@@
Let be a field and consider the dualizable category Then there exists a non-zero idempotent-complete stable subcategory such that and the intersection of with the image of is zero. In particular, the image of the homological epimorphism is not a stable subcategory.
@@
– the cohomology of the neighborhood of infinity of The first resp. second copy of corresponds to the neighborhood of resp. We denote by the corresponding idempotents in the Calkin category, and denote their images by We obtain fully faithful functors which send to resp. We claim that any non-zero object of the image of is not contained in the image of
Assume the contrary: suppose that for some sheaf the image of in the Calkin category is isomorphic to a non-zero object of Take some real number Since the constant sheaf is a compact object of it follows that the image of in is zero. Hence, the pullback of to is a compact object, i.e. for some perfect complex This holds for any hence for some Then and the image of in is isomorphic to the direct sum which is not contained in the image of This gives a contradiction.∎
@@ Appendix C Presentability of the category of dualizable categories
In this section we give an alternative, “explicit” proof of Ramzi’s theorem [Ram] which states that the category is -presentable. We also describe explicitly the -compact objects in for any uncountable regular We start with the following observation.
@@
Let be a dualizable stable category. Consider the full subcategory which consists of pairs such that the image of generates by colimits. Then the functor is an equivalence of categories.
The inverse functor is given by
@@
Indeed, for any the functor is conservative and commutes with all colimits, in particular, with geometric realizations. Hence, the functor is monadic by Lurie-Barr-Beck theorem, and we have an equivalence
It remains to check that for any the category is dualizable and the functor is strongly continuous. The latter assertion is obvious, and the former one follows from Proposition 1.56.∎
@@
Let be a presentable stable -monoidal category. Let be a regular cardinal such that is -compactly generated, the tensor product of -compact objects is -compact, and the unit object is -compact (equivalently, is an -algebra in ).
1) The category of -algebras in is -compactly generated.
2) If is uncountable, then an -algebra is -compact in if and only if the underlying object of is -compact in
@@
1) For an object denote by the tensor algebra of i.e. the free -algebra generated by Then is the left adjoint to the forgetful functor. If then our assumptions imply that Since the forgetful functor commutes with -filtered colimits, the objects are -compact in Since the forgetful functor is conservative and is -compactly generated, we deduce that the category is -compactly generated. More precisely, the category is the smallest full subcategory of which contains the objects and is closed under -small colimits.
2) The forgetful functor is monadic, and the corresponding monad on is given by the functor Since preserves -compact objects, the assertion follows from [BSY22, Lemma 6.17].∎
@@
Put
1) The category of dualizable categories is -presentable.
2) Let be an uncountable cardinal and a dualizable category. The following are equivalent.
@@
This is essentially a standard statement, we include the proof for completeness.
. Since the functors and are -strongly continuous, so is the composition
This composition is exactly the functor
. Since the functors and are -strongly continuous, so is the composition
@@
1) If is -strongly continuous, then so is
2) If is -strongly continuous, then so is
3) If and are -strongly continuous, then so is
4) If and are -strongly continuous, then so is
@@
hence it is -strongly continuous.
2) Similarly, the functor is the composition of -strongly continuous functors
hence it is -strongly continuous.
3) It suffices to prove that is in We have an exact triangle
@@
The latter functor is -strongly continuous by our assumptions. Applying Lemma C.7 (to the categories and ) we conclude that is -compact.
4) This is proved by reversing the proof of 3).∎
@@
The latter composition is -strongly continuous since both functors are -strongly continuous.
Next, we show that the functor is -strongly continuous, i.e. the object is -compact. Denote by the continuous monad on associated with Then is a monad on and we have Since is a -compact monad, the -compactness of is equivalent to the -compactness of in The latter object is simply under the identification Buy our assumptions, is -compact. ∎
@@
We first show the following.
Claim. For any dualizable category there exists a set and a monad such that
@@
Next, we prove some implications from part 3).
. Recall that is a -compact object in Hence, any dualizable object in is also -compact. This proves the implication.
. Let be a -directed system in and We only need to prove the following: if is some -strongly continuous functor such that the composition is strongly continuous, then there exists some such that the composition is -strongly continuous. To see this, note that the composition
@@
is zero. This exactly means that the composition is strongly continuous.
. Note that it is sufficient to prove that if is -compact in then the functors and for are -strongly continuous. Indeed, assuming this, we can apply Lemma C.8.
So let be in It follows directly from Proposition 1.6 that the functor is -strongly continuous: this exactly means that the spectra are -compact. To show that (the diagonal -bimodule) is -compact, choose a -small ind-system of compact (finitely presented) objects of such that By a straightforward generalization of [TV07, Proposition 2.14], for each the object is compact in Hence, the object is -compact: it is a -small filtered colimit of compact objects, namely
This proves the implication.
. First, we have a short exact sequence
Hence, satisfies and by the above it also satisfies If is a -small set, then also satisfies If is -compact, then by Lemma C.9 we see that the category satisfies
We now prove 1). By the above, it suffices to prove that the category is generated by colimits by the categories of the form where is an -compact monad. Let be an arbitrary dualizable category. Choose a generating collection of -compact objects of such that for each the object is a formal colimit of a sequence of objects of Then is a directed union of countable subsets with the same property. Denoting by the full subcategory generated by by colimits, we see that is dualizable, and the inclusion is strongly continuous. By Corollary 1.70, we have Hence, we may assume that is countable. Then is generated by a single -compact object. Arguing as in 1), we obtain an equivalence for some monad By Proposition C.2, we can find an -directed system of -compact monads such that Then This proves 1).
. Suppose that is -compact in We argue as in the proof of Claim. We see that is generated by a -small collection of -compact objects We obtain a monad such that Since is -compact in we deduce that is -compact in the category of dualizable categories under By Proposition C.1, this implies that is -compact.
We already proved that the conditions (i), (ii), (iii) and (v) are equivalent, and they are implied by (iv). It remains to prove the implication . So let be -compact in By we have a strongly continuous functor such that the image generates. Put Define the category by the pushout square in
Since is compactly generated, so is Since the categories and are -compact in so is It follows that the category is -compact in
Since the functor is fully faithful, so is Consider the quotient Then is compactly generated and is -compact in Therefore, is -compact in Hence, the short exact sequence
@@ Appendix D Urysohn’s lemma for dualizable categories
Note that the classical Urysohn’s lemma for compact Hausdorff spaces essentially says that the category (the opposite category of compact Hausdorff spaces) is generated by colimits by the unit interval The latter is easily seen to be an -compact object.
We will again use the equivalence
given by Proposition 5.3. Slightly abusing the terminology, we will call a functor strongly continuous if its left Kan extension to is strongly continuous. We denote the category of such functors by
The following proposition essentially implies Theorem D.1.
@@
1) Let be a dualizable category. The functor is a Verdier quotient. Its kernel is the full subcategory of ind-zero functors, i.e. of functors such that for any there exists such that the map is zero.
2) Denote by the right adjoint to Then for a strongly continuous functor we have
@@
The notation simply means that we look only at In other words, the last condition on means that the endofunctor commutes with existing colimits.
We will need the following basic fact about the sequential limits of directed colimits.
@@
where the two maps have components resp.
Note also that the poset of order-preserving maps is cofinal in the poset of all maps Using our assumptions on we obtain the isomorphisms
@@
Since we already know that the functor is essentially surjective, it suffices to prove part 2). Indeed, (D.1) implies that i.e. is a homological epimorphism. If is a function with required properties such that for all then the cofinality of implies that is an ind-zero functor Hence, the kernel of is generated by colimits by ind-zero functors, as required.
We now prove 2). Denote by the poset of functions with required properties. Then is a cofinal subposet of the poset of all monotone functions such that Since is directed, so is Consider another subposet which consists of monotone functions such that takes values in and is constant on and on each semi-open interval for Then is also cofinal in
Proof of Theorem D.1.
It suffices to prove the following: if a strongly continuous functor between dualizable categories induces an equivalence then is an equivalence.
It is clear that is essentially surjective. The fully faithfulness follows directly from Proposition D.2. Namely, for strongly continuous we have
Remark D.4.
Suppose that we know the cocompleteness of the category and the -compactness of in (both are not difficult to check). It is easy to deduce from the proof of Proposition D.2 that is generated by colimits by
The same applies to the category of compactly assembled presentable categories and strongly continuous right exact functors if we replace the poset with the poset with the usual order (the former poset is not a presentable category since it does not have an initial object).
@@ Appendix E Phantom maps, pure injectivity and Adams representability
Let be a dualizable category, and let be its homotopy category. As in Subsection 1.8, we denote by the ideal of compact maps in Recall that
@@
1) We say that a right -module is almost zero if it is annihilated by In other words, for any compact morphism in the map is zero.
2) We denote by the quotient of by the full (Serre) subcategory of almost zero modules.
The subcategory of almost zero modules is indeed a Serre subcategory since We make the following observation.
@@
The quotient functor has an exact left adjoint given by
@@
. If is locally coherent, then the category is abelian and the inclusion functor is exact. Passing to ind-completions, we get the functor which is therefore exact.
. Let be a morphism in We need to show that the kernel of in is also a compact object. Note that the maps and are isomorphisms. Since both functors and are left exact, we conclude that the map is also an isomorphism. This exactly means that as required.∎
@@
1) If is compactly generated, then we have
2) In general, the category is a Grothendieck abelian category which satisfies (AB4*) and (AB6). In particular, it is compactly assembled. The functor is exact.
@@
1) If is compactly generated, then is an ideal of maps in which factor through an object of Hence, the kernel of the (exact) restriction functor is the category of almost zero modules. Since its left adjoint is fully faithful, the assertion follows.
2) Recall that the category of almost zero modules is closed under limits and colimits. Since the axioms (AB4*) and (AB6) hold in they also hold in this is a special case of Roos’ theorem [Roo65]. It is clear that is a Grothendieck category. Finally, the functor is the composition of exact functors:
@@
Consider the functor which sends to considered as an almost -module.
@@
is almost zero. This shows that commutes with coproducts.
To show that is conservative, it suffices to check that if then But the vanishing of means that any compact morphism is zero. This implies that is zero.∎
Definition E.6.
Let be a morphism in
1) is phantom if for any compact morphism we have
2) is a pure monomorphism if the map is phantom.
3) is a pure epimorphism if the map is phantom.
@@
We refer to [Kr00, GP04] for the usual notions of phantom maps, pure monomorphisms and epimorphisms, and pure injective objects for compactly generated triangulated categories. We obtain the following description of the category of pure injective objects, which is a straightforward generalization of the corresponding result for compactly generated categories [GP04, Theorem 4.1].
@@
Let be an injective object of Consider the functor given by This is a cohomological functor (since is injective), which commutes with products. Hence, it is represented by some object Moreover, is a pure injective object of since annihilates phantom morphisms. We obtain a functor from the category of injective almost -modules to the category of pure injective objects of Note that we have a functorial isomorphism
Let now be a pure injective object, and choose a monomorphism where is injective. Then the map is a pure monomorphism. Since is pure injective, is a direct summand of Hence, is a direct summand of We conclude that is injective.
It remains to recall that is conservative, hence the map is an isomorphism for any pure injective ∎
Recall that if is an additive category with weak kernels, then a right -module is called fp-injective if for any finitely presented right -module we have
Proposition E.9.
is cohomological.
is flat as an -module.
is fp-injective as an -module.
Proof.
For an object we denote by the corresponding representable right -module.
Flatness of means that for any finitely presented module any morphism factors through a representable module. Let for some morphism in Then the map is given by an element of Note that is a weak cokernel of in Hence, the map factors through a representable module iff is in the image of the map We conclude that is flat if and only if is cohomological.
Let be a finitely presented module, for some morphism in Let Then we have an exact sequence of -modules
We say that an almost module is almost cohomological if it is isomorphic to an image of a cohomological -module. Equivalently, this means that the -module is flat.
@@
Indeed, for any object the -module is flat. ∎
We can now prove a generalization of Adams representability theorem for -compact dualizable categories.
Theorem E.11.
In particular, the abelian group of phantom maps is identified with The composition of any two phantom maps is zero.
Note that if is compactly generated, then is an -compact object of if and only if the triangulated category is countable. Thus, Theorem E.11 is a generalization of Neeman’s theorem [Nee97, Theorem 5.1, Proposition 4.11] if we restrict to triangulated categories with -enhancement.
Proof.
Essentially the theorem follows from the following statement.
Claim. For any the object has injective dimension at most
Assume that Claim holds. Take some objects and choose an injective resolution of of length
Consider the pure injective objects and the map (where the functor is defined in the proof of Proposition E.8). We obtain a (non-unique) map which becomes an isomorphism after applying Since is conservative, we obtain an exact triangle
@@
By construction of the functor we see that the sequence E.2 is exactly of the form (E.1).
In particular, we see that any phantom map factors through (and vice versa). If is another phantom map, then since is pure injective.
@@
By Theorem C.6 there exists a fully faithful strongly continuous functor where is -compact in Denote by the homotopy category of By Proposition 1.6, the category is countable, i.e. it has at most countably many isomorphism classes of objects, and the sets of morphisms are at most countable.
Since the functor preserves compact morphisms, we have a well-defined quotient functor
Its left adjoint is fully faithful. Moreover, takes almost cohomological almost modules to flat modules, because the extension of scalars functor preserves flatness. We also observe that is exact: composing with the conservative exact functor we obtain the functor which is isomorphic to the composition of exact functors
Now, applying Proposition E.9, we see that is fp-injective. Since is countable, it follows that has injective dimension at most Since is exact, the functor preserves injective objects. Since is also exact, we conclude that has injective dimension at most ∎

### v3

@@ Appendix B Essential image of a homological epimorphism
Consider a homological epimorphism where and are small stable idempotent-complete categories. We know that if the category is compactly generated, then In particular, the essential image of is a stable subcategory of In this case by Thomason’s theorem every object of of the form can be lifted to It turns out that if is not compactly generated then the image of might not be a stable subcategory of More precisely, the following result shows that in general for a dualizable category there may exist an object such that the object cannot be lifted to
@@
Let be a field and consider the dualizable category Then there exists a non-zero idempotent-complete full stable subcategory such that and the intersection of with the essential image of is zero. In particular, the essential image of the homological epimorphism is not a stable subcategory.
@@
– the cohomology of the neighborhood of infinity of The first resp. second copy of corresponds to the neighborhood of resp. We denote by the corresponding idempotent endomorphisms in the Calkin category, and denote their images by We obtain fully faithful functors which send to resp. We claim that any non-zero object of the image of is not contained in the essential image of Assume the contrary: suppose that for some sheaf the image of in the Calkin category is isomorphic to a non-zero object of Take some real number Since the constant sheaf is a compact object of it follows that the image of in is zero. Hence, the pullback of to is a compact object, i.e. for some perfect complex This holds for any hence for some Then and the image of in is isomorphic to the direct sum which is not contained in the essential image of This gives a contradiction.∎
@@ Appendix C Presentability of the category of dualizable categories
In this section we give an alternative, “explicit” proof of the absolute version of Ramzi’s theorem [Ram24a] which states that the category is -presentable. We also describe explicitly the -compact objects in for any uncountable regular cardinal We start with the following observation.
@@
Let be a dualizable stable category. Consider the full subcategory which consists of pairs such that the image of generates by colimits. Then the functor is an equivalence of categories. The inverse functor is given by
@@
Indeed, for any the functor is conservative and commutes with all colimits, in particular, with geometric realizations. Hence, the functor is monadic by Lurie-Barr-Beck theorem, and we have an equivalence It remains to check that for any the category is dualizable and the functor is strongly continuous. The latter assertion is obvious, and the former one follows from Proposition 1.55.∎
@@
Let be a presentable stable -monoidal category. Let be a regular cardinal such that is -compactly generated, the tensor product of -compact objects is -compact, and the unit object is -compact (equivalently, is an -algebra in ). 1) The category of -algebras in is -compactly generated. 2) If is uncountable, then an -algebra is -compact in if and only if the underlying object of is -compact in
@@
1) For an object denote by the tensor algebra of i.e. the free -algebra generated by Then is the left adjoint to the forgetful functor. If then our assumptions imply that Since the forgetful functor commutes with -filtered colimits, the objects are -compact in Since the forgetful functor is conservative and is -compactly generated, we deduce that the category is -compactly generated. More precisely, the category is the smallest full subcategory of which contains the objects and is closed under -small colimits. 2) The forgetful functor is monadic, and the corresponding monad on is given by the functor Since preserves -compact objects, the assertion follows from [BSY22, Lemma 6.17].∎
@@
Put 1) The category of dualizable categories is -presentable. 2) Let be an uncountable regular cardinal and a dualizable category. The following are equivalent.
@@
This is essentially a standard statement, we include the proof for completeness. . Since the functors and are -strongly continuous, so is the composition
This composition is exactly the functor . Since the functors and are -strongly continuous, so is the composition
@@
1) If is -strongly continuous, then so is 2) If is -strongly continuous, then so is 3) If and are -strongly continuous, then so is 4) If and are -strongly continuous, then so is
@@
hence it is -strongly continuous. 2) Similarly, the functor is the composition of -strongly continuous functors
hence it is -strongly continuous. 3) It suffices to prove that is in We have an exact triangle
@@
The latter functor is -strongly continuous by our assumptions. Applying Lemma C.7 (to the categories and ) we conclude that is -compact. 4) This is proved by reversing the proof of 3).∎
@@
The latter composition is -strongly continuous since both functors are -strongly continuous. Next, we show that the functor is -strongly continuous, i.e. the object is -compact. Denote by the continuous monad on associated with Then is a monad on and we have Since is a -compact monad, the -compactness of is equivalent to the -compactness of in The latter object is isomorphic to under the identification By our assumptions, is -compact. ∎
@@
We first show the following. Claim. For any dualizable category there exists a set and a monad such that
@@
Next, we prove some implications from part 2). . Recall that is a -compact object in Hence, any dualizable object in is also -compact. This proves the implication. . Let be a -directed system in and We only need to prove the following: if is some -strongly continuous functor such that the composition is strongly continuous, then there exists some such that the composition is -strongly continuous. To see this, note that the composition
@@
is zero. This exactly means that the composition is strongly continuous. . Note that it is sufficient to prove that if is -compact in then the functors and for are -strongly continuous. Indeed, assuming this, we can apply Lemma C.8. So let be in It follows directly from Proposition 1.6 that the functor is -strongly continuous: this exactly means that the spectra are -compact. To show that (the diagonal -bimodule) is -compact, choose a -small ind-system of compact (finitely presented) objects of such that and denote by the functors to the colimit. By a straightforward generalization of [TV07, Proposition 2.14], for each the object is compact in Hence, the object is -compact: it is a -small filtered colimit of compact objects, namely
This proves the implication. . First, we have a short exact sequence
from Proposition 4.21. Hence, satisfies (iv) and by the above it also satisfies (iii). If is a -small set, then also satisfies (iii). If is -compact, then by Lemma C.9 we see that the category satisfies (iii). We now prove 1). By the above, it suffices to prove that the category is generated by colimits by the categories of the form where is an -compact monad. Let be an arbitrary dualizable category. Choose a generating collection of -compact objects of such that for each the object is a formal colimit of a sequence of objects of Then is a directed union of countable subsets with the same property. Denoting by the full subcategory generated by by colimits, we see that is dualizable, and the inclusion is strongly continuous. By Corollary 1.69, we have Hence, we may assume that is countable. Then is generated by a single -compact object. Arguing as in 1), we obtain an equivalence for some monad By Proposition C.2, we can find an -directed system of -compact monads such that Then This proves 1). . Suppose that is -compact in We argue as in the proof of Claim. By the above, we see that is generated by a -small collection of -compact objects We obtain a monad such that Since is -compact in we deduce that is -compact in the category of dualizable categories under By Proposition C.1, this implies that is -compact. We already proved that the conditions (i), (ii), (iii) and (v) are equivalent, and they are implied by (iv). It remains to prove the implication . So let be -compact in By (ii), we have a strongly continuous functor such that the image generates. Put Define the category by the pushout square in
Since is compactly generated, so is Since the categories and are -compact in so is It follows that the category is -compact in Since the functor is fully faithful, so is Consider the quotient Then is compactly generated and is -compact in Therefore, is -compact in Hence, the short exact sequence
@@ Appendix D Urysohn’s lemma for dualizable categories
Note that the classical Urysohn’s lemma for compact Hausdorff spaces essentially says that the category (the opposite category of compact Hausdorff spaces) is generated by colimits by the unit interval The latter object is easily seen to be -compact in see Proposition F.5 below.
Proof of Theorem D.1.
We will consider the set as a linearly ordered set with the usual order. For we consider the subsets also as linearly ordered sets. We will again use the equivalence
given by Proposition 5.3, where is a dualizable category. By Theorem C.6 we already know that the category is presentable and the object is -compact. Thus, we only need to prove the following: if a strongly continuous functor between dualizable categories induces an equivalence
then is an equivalence. We first recall that the functor
is essentially surjective by Corollary C.5, hence is also essentially surjective. To prove the fully faithfulness of we need the following observation: for any pair of strongly continuous functors we have pullback squares of spaces:
and
Since the posets and are isomorphic to the poset we conclude that the map
is an equivalence of spaces. Now take any objects and choose strongly continuous functors such that Then we obtain the following chain of equivalences of spaces, where all the limits and colimits are over real numbers:
Here the second equivalence follows from the isomorphism The third equivalence follows from the observation that for any we have an ind-equivalence The fourth map is an equivalence since (D.2) is an equivalence and we have obvious equivalences of posets Now, a similar chain of equivalences identifies the target of (D.3) with This proves the fully faithfulness of and the theorem. ∎
Remark D.2.
Suppose that we know the cocompleteness of the category and the -compactness of the poset in (both are not difficult to check). Then the same argument as in the proof of Theorem D.1 shows that the category is generated via colimits by the poset A similar statement holds for the category of compactly assembled presentable categories and strongly continuous right exact functors. In this case the -compact generating object is the category
For completeness we explain a direct way to recover the dualizable category from the category of strongly continuous functors Slightly abusing the terminology, we will call a functor strongly continuous if its left Kan extension to is strongly continuous. We denote the category of such functors by
@@
1) Let be a dualizable category. The functor
is a Verdier quotient. Its kernel is the full subcategory of ind-zero functors, i.e. of functors such that for any there exists such that the map is zero. 2) Denote by the right adjoint to Then for a strongly continuous functor we have
@@
The notation simply means that we look only at In other words, the last condition on means that the endofunctor commutes with existing colimits. We will need the following basic fact about sequential limits of directed colimits.
@@
where the two maps have components resp. Note also that the poset of order-preserving maps is cofinal in the poset of all maps Using our assumptions on we obtain the isomorphisms
@@
We already know that the functor is essentially surjective by Corollary C.5. Hence, it suffices to prove part 2). Indeed, (D.4) implies that i.e. is a homological epimorphism. Let be a function with required properties such that for all Then for any strongly continuous functor the functor is an ind-zero functor (since is cofinal). Hence, the kernel of is generated by colimits by ind-zero functors, as required. We now prove 2). Denote by the poset of functions with required properties. Then is a cofinal subposet of the poset of all (not necessarily strictly) monotone functions such that Since is directed, so is Consider another subposet which consists of monotone functions such that takes values in and is constant on and on each semi-open interval for Then is also cofinal in Consider two strongly continuous functors Note that for the functor is right Kan-extended from Applying Lemma D.4 and using the isomorphism we obtain the isomorphisms
This proves (D.4). ∎
@@ Appendix E Phantom maps, pure injectivity and Adams representability
In this section we prove two closely related analogues of Adams representability theorem for -compact dualizable categories (Theorems E.2 and E.4). Let be a dualizable category, and denote by the triangulated homotopy categories of and We say that a morphism in is compact if it is an image (i.e. homotopy class) of a compact morphism in The following are straightforward generalizations of the usual notions for compactly generated categories [Kr00, GP04].
@@
Let be a morphism in 1) is phantom if for any compact morphism with we have 2) is a pure monomorphism if the map is phantom. 3) is a pure epimorphism if the map is phantom.
Recall that for a small triangulated category an additive functor is called homological if for any exact triangle in the sequence is exact. Clearly, homological functors are closed under extensions in the abelian category of left -modules (i.e. the category of additive functors ). We will also refer to homological functors as cohomological.
Theorem E.2.
Let be a dualizable category which is is -compact in We denote by resp. the triangulated homotopy category of resp. Consider the functor given by Then the functor is full and conservative, and its essential image consists of homological functors which commute with countable coproducts. Moreover, for we have a functorial short exact sequence of abelian groups
The subgroup of phantom maps is identified with
We will deduce Theorem E.2 from the closely related Theorem E.4 below. To formulate it, we need some preparations. We denote by resp. the ideal of compact maps in resp. Recall from Subsection 1.8 that
Definition E.3.
1) We say that a right -module is almost zero if it is annihilated by In other words, for any compact morphism in the map is zero. The definition of an almost zero left -module is similar. 2) We denote by the quotient of by the full (Serre) subcategory of almost zero right modules. Similarly, we denote by the quotient of by the full subcategory of almost zero left modules.
The subcategories of almost zero modules are indeed Serre subcategories since and If is compactly generated, then both categories and are identified with (Proposition E.9 below). We say that an almost module is almost cohomological if it is an image of a right -module corresponding to a cohomological functor
Theorem E.4.
The abelian group of phantom maps is identified with The composition of two phantom maps is zero.
Note that if is compactly generated, then is an -compact object of if and only if the triangulated category is countable. Thus, Theorem E.4 is a generalization of Neeman’s theorem [Nee97, Theorem 5.1, Proposition 4.11] if we restrict to triangulated categories with an -enhancement. Before proving Theorem E.4 we explain some basic properties of the abelian category and show that it is naturally equivalent to First we recall the following characterization of (co)homological functors. Recall that if is a small additive category with weak kernels, then a right -module is called fp-injective if for any finitely presented right -module we have
@@
is cohomological.
is flat as a right -module.
is fp-injective as a right -module.
@@
For an object we denote by the corresponding representable right -module. Flatness of means that for any finitely presented module any morphism factors through a representable module. Let for some morphism in Then the map is given by an element Note that is a weak cokernel of in Hence, the map factors through a representable module iff is in the image of the map We conclude that is flat if and only if is cohomological. Let be a finitely presented module, for some morphism in Let Then we have an exact sequence of -modules
Next, recall from Proposition 1.48 that the tensor product is the quasi-ideal in given by We denote by the similar quasi-ideal in given by We make the following observation.
Proposition E.6.
The quotient functor resp. has an exact left adjoint resp. given by
resp.
Proof.
Proposition E.7.
Consider the colimit-preserving functor which is given on representable modules by Then is exact and it vanishes on almost zero right -modules. The induced functor
is an equivalence.
Proof.
Exactness of means that for an object the left -module is flat. This follows from Proposition E.5. By Proposition E.6, to show the vanishing of on almost zero modules, we need to check that for any the map is an isomorphism. By exactness of and we may assume for some Let (in ). Then and
This proves the vanishing of on almost zero modules. Now recall the functor from Remark 1.12 (with the roles of and interchanged). To avoid confusion, we denote the corepresentable left -modules by Consider the colimit-preserving functor given on corepresentable objects by We claim that Indeed, let with Choose factorizations in with Then we have and
The isomorphism is functorial in hence we get an isomorphism as stated. Arguing as above, we see that the functor is exact. Now to prove that the functor (E.2) is an equivalence, it suffices to construct the isomorphisms
We only need to construct these isomorphisms functorially on corepresentable left -modules. So let and let in Choose a pro-system in such that Then and we obtain
Similarly, we obtain
This proves the proposition. ∎
@@
. If is locally coherent, then the category is abelian and the inclusion functor is exact. Passing to ind-completions, we get the functor which is therefore exact. . Let be a morphism in We need to show that the kernel of in is also a compact object. Note that the maps and are isomorphisms. Since both functors and are left exact, we conclude that the map is also an isomorphism. This exactly means that as required.∎
@@
1) If is compactly generated, then we have 2) In general, the category is a Grothendieck abelian category which satisfies (AB4*) and (AB6). In particular, it is compactly assembled. The functor is exact.
@@
1) If is compactly generated, then is an ideal of maps in which factor through an object of Hence, the kernel of the (exact) restriction functor is the category of almost zero modules. Since its left adjoint is fully faithful, the assertion follows. 2) Recall that the category of almost zero modules is closed under limits and colimits. Since the axioms (AB4*) and (AB6) hold in they also hold in this is a special case of Roos’ theorem [Roo65]. It is clear that is a Grothendieck category. Finally, the functor is the composition of exact functors:
@@
Consider the functor from Theorem E.4, i.e. considered as an almost -module.
@@
is almost zero. This shows that commutes with coproducts. To show that is conservative, it suffices to check that if then But the vanishing of means that any compact morphism is zero. This implies that is zero.∎
@@
We refer to [Kr00, GP04] for the usual notion of pure injective objects for compactly generated triangulated categories. We obtain the following description of the category of pure injective objects, which is a straightforward generalization of the corresponding result for compactly generated categories [GP04, Theorem 4.1].
@@
Let be an injective object of Consider the functor given by This is a cohomological functor (since is injective), which commutes with products. Hence, it is represented by some object Moreover, is a pure injective object of since annihilates phantom morphisms. We obtain a functor from the category of injective almost -modules to the category of pure injective objects of Note that we have a functorial isomorphism Let now be a pure injective object, and choose a monomorphism where is injective. Then the map is a pure monomorphism. Since is pure injective, it follows that is a direct summand of Hence, is a direct summand of We conclude that is injective. It remains to recall that is conservative, hence the map is an isomorphism for any pure injective ∎
@@
Indeed, for any object the -module is cohomological. ∎
Proof of Theorem E.4.
Essentially the theorem follows from the following statement. Claim. Any almost cohomological almost -module has injective dimension at most Assume that Claim holds. Let be an almost cohomological almost -module. Choose an injective resolution of of length
Consider the pure injective objects and the map where the functor is defined in the proof of Proposition E.12. Then the map is a pure epimorphism. Putting and using the fact that is cohomological, we obtain
Hence, the essential image of is exactly the full subcategory of almost cohomological almost modules. Now take some objects and choose an injective resolution of as above:
We obtain a (non-unique) map which becomes an isomorphism after applying Since is conservative, we obtain an exact triangle
@@
By construction of the functor we see that the sequence (E.3) is exactly of the form (E.1). In particular, we see that any phantom map factors through (and vice versa). If is another phantom map, then since is pure injective.
@@
By Theorem C.6 there exists a fully faithful strongly continuous functor where is -compact in Denote by the homotopy category of By Proposition 1.6, the category is countable, i.e. it has at most countably many isomorphism classes of objects, and the sets of morphisms are at most countable. Since the functor preserves compact morphisms, we have a well-defined quotient functor
Its left adjoint is fully faithful. Moreover, takes almost cohomological almost modules to flat modules, because the extension of scalars functor preserves flatness. We also observe that is exact: composing with the conservative exact functor we obtain a functor which is isomorphic to the composition of exact functors
where as above Now, applying Proposition E.5, we see that for any almost cohomological almost module the -module is fp-injective. Since is countable, it follows that has injective dimension at most Since is exact, the functor preserves injective objects. Since is also exact, we conclude that has injective dimension at most ∎
@@
Proof of Theorem E.2.
The proof is obtained from Theorem E.4 using the relation between left -modules and right -modules. We say that a left almost module is almost homological if it is an image of a left -module corresponding to a homological functor Recall from Proposition E.6 that the (exact, fully faithful) left adjoint to the quotient functor is given by where is the quasi-ideal in given by The essential image of the functor consists of functors such that the composition commutes with sequential colimits. It follows that the functor gives an equivalence
Indeed, this follows from the observation that for a homological functor the composition commutes with sequential colimits if and only if commutes with countable coproducts. Therefore, in the formulation of the theorem we can replace the functor with the composition
and prove that this functor has the same properties as the functor from Theorem E.4. It remains to observe that the following square commutes
where is the equivalence from Proposition E.7. It is clear that sends almost homological left almost -modules to almost cohomological right almost -modules. Therefore, the theorem follows from Theorem E.4. ∎